# Image Matching Filtering and Refinement by Planes and Beyond

### v3

@@ 4.2.1 Prelude
Tables 1-7 report the results obtained according to the setup presented in Sec. 4.1 for the different base pipelines. The left aligned shift in the pipeline column indicates incremental inclusions of the modules as in the folder-tree style visualization. All the values in the tables are expressed in percentages. Extended tables including AUCs values in the non-aggregated form, i.e. for the different thresholds, are reported as online resources6. In any case, non-aggregated AUC ranks are almost stable and well correlated with the corresponding aggregated measure.
@@
In the next discussion, the base pipelines will be referred to for brevity by the first component, e.g. Key.Net will refer to Key.Net++NNR. CC filtering was not evaluated on IMC-PT since it experienced a high number of unreasonable crashes for this dataset.
Referring to the tables, pose estimation from the fundamental matrix is harder than that from the essential matrix. This can be observed by the value differences in the sided green and yellow columns. Moreover, using MAGSAC before RANSAC-based essential matrix estimation in general does not affect or even slightly decreases the accuracy. According to these observations, the behavior of the different matching filters is stressed when the pose is estimated from the fundamental matrix. For IMC-PT the corresponding angular (green and yellow columns) and metric (purple and olive green columns) AUC scores maintain the same relative order even if metric scores are lower in terms of absolute value, thus highlighting that real pose solution is still far to be solved.
The previous observations are further confirmed by the analysis of the correlation within the employed error measures, reported in Table 8 for each dataset. On one hand, after the MAGSAC geometric constraints, for both planar and non-planar scenes there is a high correlation within recall, pose estimation accuracy as AUC, regardless of the estimation is derived from the fundamental or essential matrix, respectively indicated by the ‘’ and ‘’ superscripts, and these both are highly anti-correlated with the percentage of filtered matches. On the other hand, without MAGSAC in the non-planar sceness the recall and the pose accuracy estimated by the essential matrix are highly correlated and both strongly anti-correlated with the number of filtered matches, while precision has an almost moderate correlation with the pose accuracy estimated from the fundamental matrix only. Metric and angular AUC for the same kind of pose estimation have also high correlation levels. For the planar case, without MAGSAC the pose accuracy is only moderately correlated with both precision and recall, as well as anti-correlated with the percentages of the filtered matches within a similar degree.
For the sake of completeness, Table 9 reports the average number of input matches for each base method and dataset under evaluation. This information can be used together with the percentage of filtered matches to get an absolute estimation of the filtered matches. According to the reported values, SuperPoint and ALIKED provide the lower number of matches followed by SIFT and Key.Net, while DISK, LoFTR and DeDoDe v2 the highest number.
@@ 4.2.2 SIFT+NNR
Table 1 reports the results for SIFT. MOP+MiHo obtains the best pose estimation on Megadepth when this is computed from the essential matrix (AUC, yellow column), while adding NCC slightly decreases the accuracy. MOP-based filters, FC-GNN, and ACNe pose estimation are more accurate than the base SIFT even if the improvement is incremental. Other matching filters yield minimal degradation of the base accuracy in this specific setup. When the pose is derived from the fundamental matrix (AUC, green column), FC-GNN attains the highest score followed by MOP, which is almost unaltered by MiHo and again negatively altered by NCC. The improvement with respect to the base SIFT is more evident for AUC than for AUC. All filters generally improve on SIFT considering AUC, but only MOP-base filters, FC-GG, ACNe, and CC are really remarkable in terms of the accuracy gain. This behavior is also reflected in the recall (red column) and precision (blue column). In particular, MOP+MiHo seems to provide a better recall and FC-GNN more precision, where the recall is more correlated with AUC and the precision with AUC as discussed in Sec. 4.2.1.
Similar considerations hold for ScanNet even if the absolute scores are lower due to a major scene complexity, yet NCC improves MOP+MiHo and ACNe accuracy is relatively closer to those of MOP-based filters and FC-GNN. Specifically, ACNe achieves the best AUC score followed by MOP+MiHo+NCC and then FC-GNN, while FC-GNN and ACNe are marginally better than MOP-based filters in terms of AUC. Moreover, ConvMatch, DeMatch, and OANet are better in pose accuracy than the base SIFT.
CC and MOP+MiHo+NCC obtain the best pose accuracy (AUC, purple column) in planar scenes as they both rely on homographies. The base MOP is better in this case than MOP+MiHo and MOP+MiHo+NCC is better than base MOP. Other filters cannot overstep the base MAGSAC in the planar scenario. Moreover, the most aggressive filters in terms of filtered matches (gray columns), i.e. AdaLAM, OANet, and MS2DG-Net, are those with worse results. Since the NNR threshold was set deliberately high, this suggests a lack of adaptability to the input matches.
Finally, all filters provide better AUC score than the base SIFT in the non-planar IMC-PT dataset, with the exception of AdaLAM, MS2DG-Net, and GMS, but without noticeable differences as already seen for MegaDepth. ACNe achieves the best score, while FC-GNN improvements are smaller compared to those obtained for MegaDepth. For AUC the ranking is quite similar to that obtained for of ScanNet, with the difference that CLNet and NCMNet are better than the base SIFT. Analogously observations can be done for the metric pose accuracy (AUC and AUC, respectively purple and olive green columns) but, in this case, ACNe is preceded by MOP-base filtering and FC-GNN. Additionally, NCC improves MOP+MiHo for the images in this dataset.
@@
As previously mentioned and confirmed by a private discussion with authors, ACNe has a possible and consistent overlap with the training data used for the estimation of the indoor and outdoor model weights with ScanNet and IMC-PT, respectively. Moreover, although training data used to compute FC-GNN weights and the validation data are different for Megadepth, the kinds of scenes are similar. For SIFT as well as other base pipelines discussed next, the gap within AUC accuracy computed with fundamental and essential matrix estimations of FG-CNN is quite limited for MegaDepth and increases in order with IMC-PT and ScanNet. This suggests, together with the analysis of the intrinsic camera parameters reported in Fig. 5, the possibility that FC-GGN implicitly learned the camera intrinsic priors for the specific MegaDepth dataset, partially overlapping with IMC-PT, that cannot be always exploited in other non-planar datasets. Differently, MOP is not tuned on the data being handcrafted.
@@
Table 2 reports the the results for Key.Net. Differently from SIFT, only MOP-based filtering and FC-GNN achieve better pose accuracy than the base Key.Net in the case of AUC and AUC, respectively, in the all non-planar datasets with or without MAGSAC. AUC score is higher for MOP+MiHo+NCC than FC-GNN and vice-versa in the case of AUC. This consideration also holds for the metric scores AUC and AUC. CC and MOP-based filter provide the best AUC scores in the planar dataset as for SIFT.
Overall, the best score achieved by any Key.Net pipeline is lower than the corresponding SIFT one with the relevant exception of Key.Net with MOP-based filtering and NCC. For this pipeline, Key.Net plus MOP+MiHo+NCC obtains the best absolute scores also among SIFT-based pipelines when pose accuracy is estimated from the essential matrix. This indirectly supports the different nature of the SIFT blobs and Key.Net corners. From the authors’ experience, on one hand, blobs are characterized to be more flat-symmetric and hence less discriminative than corners. However, on the other hand, for the same reasons keypoint localization is more stable for blobs than corners. A proper refinement step as NCC thus improves the matching process for corner-like features. Notice also that the recall improvement obtained by adding NCC with respect to the same pipeline without NCC, and in minor part also the precision, give indications of previous matches that were outliers due to a low keypoint accuracy and have been turned into inliers.
Moreover, filters yield worse scores than the base pipeline plus MAGSAC for Key.Net, with the exception of FC-GNN and MOP-based filtering. This highlights a critical aspect of learning-based methods, that need an appropriate tuning and training on the base of the other pipeline components. Notice also that the difference between the corresponding AUC and AUC for FC-GNN on MegaDepth is more limited than with other datasets, and in particular with respect to IMC-PT. A possible justification is that the network is able to learn camera intrinsic priors of the MegaDepth images, which have less variability as outdoor scenes than IMC-PT. As discussed in the evaluation setup, all deep filters have been trained on SIFT, including FC-GNN, suggesting less generalization to different image features with respect to handcrafted approaches based on geometric constraints like MOP.
Finally, unlike SIFT, directly applying NCC after the base pipeline without MOP provides some improvements, even if these are inferior to when NCC is preceded by MOP-base filtering, confirming the original assumption that affine patch normalization is better than normalization by similarity for the refinement, but not as good as planar patch normalization.
@@ 4.2.4 SuperPoint+LightGlue
Table 3 reports the results for SuperPoint. Overall, the relative rank among the datasets and pipelines for SuperPoint is comparable with Key.Net. This is not accidental since both Key.Net and SuperPoint are based on corner-like keypoints so the same considerations and observations done for Key.Net hold for SuperPoint. Moreover, in terms of absolute score, SuperPoint with MOP+MiHo+NCC achieves on MegaDepth the top AUC value among all the evaluated pipelines, or a value almost equal to the best absolute AUC provided by LoFTR for ScanNet. Finally, the improvement in adding NCC in terms of recall and precision with respect to Key.Net is inferior but still relevant, implying a more accurate base corner relocalization of SuperPoint.
4.2.5 ALIKED+LightGlue
Table 4 reports the results for ALIKED. Overall, no method with or without MAGSAC has better AUCs than the base ALIKED with or without MAGSAC, respectively. All filters strongly degrade AUC scores with the exception of MOP-based filtering, FC-GNN, and AdaLAM. AdaLAM also has exactly the same AUCs as the base ALIKED. This can be interpreted positively since in case of perfect input matches a filters should not alter it. FC-GNN provides better results than the base ALIKED only on MegaDepth, while MOP+MiHo with or without NCC provides more or less the same AUCs of the base pipeline in the non-planar cases. As for AdaLAM, this is a positive feature. In the planar dataset, only MiHo-based filtering is remarkably better than the base pipeline while most of the other filters are slightly inferior to the base pipeline. Note also that corresponding SuperPoint pipelines are better than the ALIKE ones in terms of the best absolute AUC scores.
ALIKED provides overall more robust and accurate base input matches than SuperPoint, also in terms of keypoint localization. As a downside, this strict accuracy of the input limits the number of less probable correct matches which could be better filtered or adjusted by a robust filtering approach.
4.2.6 DISK+LightGlue
Table 5 reports the results for DISK. The relative rank is generally quite similar to that obtained by ALIKED for the non-planar scenes. The main difference is that FC-GNN is marginally better than the base DISK for any dataset and not only for MegaDepth, while MOP-based filtering behaves as the base pipeline and NCC always degrades slightly the AUC. This is reasonable considering that DISK keypoints are more similar to blobs than corners as one can note by visual inspection. Finally, in terms of absolute scores, the top-rank AUCs for IMC-PT are better than those on SuperPoint but much lower on ScanNet.
Concerning the planar scenes, MAGSAC decreases the final pose accuracy and, with the exception of NCMNet and FC-GNN, all other filters, including MOP-based ones, are noticeably worse than the base pipeline in terms of AUC. In particular, CC is the filter with lower AUC.
@@ 4.2.7 LoFTR
Table 6 reports the results for LoFTR. Unlike other architectures relying on SuperGlue, MAGSAC greatly improves LoFTR base pipeline. Similarly to ALIKED, except AdaLAM, FC-GNN, and MOP-based filtering, other filters are below the base LoFTR with MAGSAC in non-planar scenes. In particular, AdaLAM, MOP, and MOP+MiHo are equal to LoFTR with MAGSAC in outdoor datasets, yet the refinement with FC-GNN or by MOP-based filtering and NCC greatly improves AUCs. This is highlighted by the high precision and the recall values exceeding those of the base LoFTR when using MOP-based filtering with NCC and without MAGSAC. As for previous base pipelines, FC-GNN is better than MOP-based filters with NCC in the case of pose estimation from the fundamental matrix. Nevertheless, refinement by NCC or FC-GNN introduces score losses in the indoor ScanNet. In terms of absolute score, top AUC scores in the case of estimation of the pose by the fundamental matrix are higher than the corresponding one of SuperGlue in the outdoor scenes.
For planar images, homography estimation accuracy in terms of AUC is improved in order for CC, CLNet, NCMNet, and FC-GNN, but MOP-based NCC refinements achieve the top scores, which are also the best as absolute values with respect to other pipelines. Notice that MOP and MOP+MiHo without NCC achieve good performances in between those previously mentioned.
4.2.8 DeDoDe v2
Table 7 reports the results for DeDoDe v2. MAGSAC improves pose accuracy according to AUC or AUC, and AdaLAM has the same scores of the base DeDoDe v2 for every dataset. Moreover, both CC and GMS in general have better AUC scores than the base pipeline in outdoor datasets. MOP and MOP+MiHo give improvements excluding the indoor datasets where the original score is maintained, and NCC inclusion degrades the accuracy for the outdoor datasets while improving it for indoor and planar scenes. Excluding the planar scenes, FC-GNN achieved the best results even if strictly better than MOP only for AUC, being comparable to MOP-based filtering in the case of AUC.
4.2.9 Visual qualitative analysis
Figure 7 shows visual results of patch alignment by NCC for two example image pairs, superimposed as red-green anaglyphs. Patches have been randomly chosen for visualization, and full patch sets for each evaluated pipeline and dataset combination can be generated by the provided supporting tool1. Inside each multirow in the figure, the top row shows the initial patch alignment specified by the homography pair as defined in Sec. 3.3, while the second row shows the results after applying NCC without MOP-based homography solutions. Likewise, the third row shows the initial alignment of the MOP+MiHo extended homography pair and the bottom row the final results with NCC. In the left clusters , with or without NCC, has the highest correlation value. The blue bar on the left of each patch superimposition indicates the corresponding epipolar error of the patch center with respect to the pseudo GT. The reader is invited to zoom in and inspect the above figure, which provides examples of the possible configurations and alignment failure cases.
For true planar regions, such as the patches including the St. Peter’s Basilica epigraph, it is evident that MOP+MiHo extended patches are better than the base ones in terms of alignment and distortions with respect to the original images. Nevertheless, MOP patch alignment can go wrong in some circumstances as shown in the leftmost column of the top multirow. This happens for instance when the homography assignment described in Sec. 3.1.4 picks a homography with a small reprojection error but also a poor wrong support set. Furthermore, NCC similarity maximization is not always the best choice as highlighted by the patches containing the inscription in the left clusters, where the registration is visibly better with MOP although base patches achieve the highest NCC correlation value.
In the case of patches representing almost flat regions or straight contour edges according to the Harris corner definition [62], NCC can be misled and degrade the alignment increasing the epipolar error, as in the rightmost patch overlay in the third multirow. Notice also that although visually improving the overall patch alignment, NCC can increase the epipolar error of the effective keypoint center, as for the ‘BV’ inscription in the third multirow. Anyways, when NCC keypoint accuracy decreases assuming that the pseudo-GT sub-pixel estimation can be always trusted, the epipolar error increment is generally limited as reported in the figure and by the results of Tables 1-7.
Figures 8-9 present some example image pairs, respectively in the case of non-planar and planar scenes, for base matching pipelines where NCC refinement is effective according to the quantitative evaluation. Base matches after being filtered by MOP+MiHo+NCC and FC-GNN, with or without RANSAC, are reported, where the match color indicates the corresponding maximum epipolar error within the pair images.
Comparing the first row for each scene, i.e. without RANSAC-based filtering, both MOP+MiHo and FC-GNN remove outliers, highlighted in blue, but MOP+MiHo is more relaxed in the task, acting as a sort of coarse Local Optimized RANSAC (LO-RANSAC) [109] step. As reported in Tables 1-7, this is beneficial for the effective RANSAC especially when the outlier contamination of the input matches is high, such as for SIFT and Key.Net, and does not negatively affect the final results in the other cases. Concerning the refinement, both NCC and FC-GNN are able to improve the keypoint accuracy, as indicated by the color shift of the matches, but the epipolar error variation of their refined matches differ. For NCC the same observations and considerations discussed for Figure 7 hold.

### v4

@@ 4.2.1 Prelude
In the next discussion, the base pipelines will be referred to for brevity by the first component, e.g. Key.Net will refer to Key.Net++NNR. Tables 1-5 report the results obtained according to the setup presented in Sec. 4.1 for SIFT, Key.Net, SuperPoint, LoFTR and ALIKED. For DISK and DeDoDe v2 results are similar to ALIKED and are reported in Appendix E. The left aligned shift in the pipeline column indicates incremental inclusion of the modules as in the folder-tree style visualization. All the values in the tables are expressed in percentages. Extended tables including AUCs values in the non-aggregated form, i.e. for the different thresholds, are provided as online material6. In any case, non-aggregated AUC ranks are almost stable and well correlated with the corresponding aggregated measure.
@@
CC filtering was not evaluated on IMC-PT because it experienced a high number of unreasonable crashes for this dataset, neither Progressive-X and Keypt2Subpx, since they were included later and according to the preliminary analysis further evaluations would have not provided any interesting additions.
There are obvious factors to take into account when analyzing the results:
When using a robust outlier estimation (i.e. MAGSAC or the base essential matrix estimation which implicitly runs RANSAC) small differences in AUC roughly within 0.5%, if not exhibiting regularly among the different pipelines and datasets, are not effectively relevant since mainly due to randomization.
The estimation of the fundamental and homography matrices without a robust inlier framework becomes basically a least-square problem so that a few strong outliers may affect the final result very badly.
A low number of very accurate matches may be not sufficient for a precise fundamental matrix estimation, not only due to possible degenerate configurations (i.e. matches only lying on a plane), but also because epipolar line estimations far from the point locations of the matches (i.e. outside the related convex hulls in the images) generally become inaccurate.
That said, having more matches on the images, even if less accurate but better distributed, can lead to a better fundamental matrix estimation. It depends on the specific input images and matches, the pipeline, the RANSAC employed and its parameters.
More constrained problems, i.e. homography and essential matrix estimations, are more robust to less accurate matches.
According to above points, referring to the tables, the pose estimation from the fundamental matrix is harder than that from the essential matrix. This can be observed by the value differences in the sided green and yellow columns. Moreover, using MAGSAC before RANSAC-based essential matrix estimation in general does not affect or even slightly decreases the accuracy. The behavior of the different matching filters is stressed when the pose is estimated from the fundamental matrix. For IMC-PT the corresponding angular (green and yellow columns) and metric (purple and olive green columns) AUC scores maintain the same relative order even if the metric scores are lower in terms of absolute values, thus highlighting that real pose is still far to be found. The previous observations are further confirmed by the analysis of the correlation between the employed error measures, reported in Appendix E.1. The average number of input matches for each base method and dataset is provided in Appendix E.2.
@@
According to the previous observations, the score achieved by filters without MAGSAC will be taken into consideration for the essential matrix estimation and results obtained with MAGSAC in the case of fundamental matrix or homography estimations.
The detailed comparative analysis of the considered filtering modules for the different base matching pipelines in terms of pose accuracy, as well as the recall and precision of the matches, is reported hereafter.
@@ 4.2.2 SIFT+NNR
Table 1 reports the results for SIFT. MOP+MiHo works overall quite well, achieving top rank scores or close to them in all dataset and configurations. Adding NCC somewhat degrades the scores for the fundamental matrix estimation with MAGSAC on MegaDepth and IMC-PT, probably due to a prominent presence of flat blob-like keypoint patches which are not suitable for NCC template matching. FC-GNN obtains top rank scores or close to these, while ACNe is quite good for ScanNet and both IMC-PT and Progressive-X on MegaDepth. Most of the other filters do better than the base pipeline for the non-planar datasets, but only for those mentioned above the gain is remarkable.
In the case of planar scenes, the differences of the filtering pipelines with respect to the base pipeline are quite limited. Filters based on planes like the MOP-based ones, CC and Progressive-X are generally more stable, in the sense that they do not degrade the base solution. Notice than FG-CNN is effectively worse than the base pipeline with the planar scenes.
MiHo reduces the MOP overfit towards planar structures, a behavior that the base MOP has in common with CC and Progressive-X due to their designs. This is highlighted by the higher number of filtered matches of MOP without MiHo that leads in turn to a higher AUC accuracy without MAGSAC for the fundamental or homography estimations. Nevertheless, including robust estimators in the pipeline inverts or cancels out this trend as discussed in Sec. 4.2.1.
As final consideration, FC-GNN and MOP-based filters tend to reach the highest level of precision and recall possible, respectively. This in turn leads to have high accuracy in term of AUC for fundamental, essential matrix estimations for FC-GNN and MOP-based filters, respectively.
@@
As previously mentioned and confirmed by a private discussion with authors, ACNe has an overlap with the training data used for the estimation of the indoor and outdoor model weights with ScanNet and IMC-PT, respectively. Moreover, although training data used to compute FC-GNN weights and the validation data are different for Megadepth, the kinds of scenes are similar. For SIFT as well as other base pipelines discussed next, the gap within AUC accuracy computed with fundamental and essential matrix estimations for FG-CNN is quite limited in the case of MegaDepth and increases in order with IMC-PT and ScanNet. This suggests, together with the analysis of the intrinsic camera parameters reported in Appendix. D.2, the possibility that FC-GGN implicitly learned the camera intrinsic priors for the specific MegaDepth dataset, partially overlapping with IMC-PT, that cannot be always exploited in other non-planar datasets. Differently, being handcrafted, MOP is not tuned on the data.
@@
Table 2 reports the the results for Key.Net. The best scores for all datasets are obtained by FC-GNN and MOP+MiHo+NCC. The inclusion of NCC in this case is effective, since the kind of keypoints provided by the base Key.Net pipeline are almost corners by design, where NCC template matching is robust. As for SIFT, FC-GNN tends to increase more the precision while MOP-based filters the recall, leading to the previous observations. Again, FC-GNN worsens the performances of the base pipeline with MAGSAC on the planar dataset.
Overall, the best score achieved by any Key.Net pipeline with or without matching filters is lower than the corresponding SIFT one, with the relevant exception of Key.Net with MOP-based filtering and NCC. Within the MOP-based pipelines, Key.Net plus MOP+MiHo+NCC obtains the best absolute scores also among SIFT-based pipelines when the pose accuracy is estimated from the essential matrix. This indirectly supports the different nature of the SIFT blobs and Key.Net corners. From the authors’ experience, on one hand, blobs are characterized to be more flat-symmetric and hence less discriminative than corners. However, on the other hand,for the same reasons keypoint localization is more stable for blobs than corners. A proper refinement step as NCC thus improves the matching process for corner-like features. Notice also that the recall improvement obtained by adding NCC with respect to the same pipeline without NCC, and in minor part also the precision, give indications of previous matches that were outliers due to a low keypoint accuracy and have been turned into inliers.
All the filters yield worse scores than the base pipeline plus MAGSAC for Key.Net, with the exception of FC-GNN and MOP-based ones. This highlights a critical aspect of learning-based methods, that need an appropriate tuning and training according to the specific pipeline components. Notice also that the difference between the corresponding AUC and AUC for FC-GNN on MegaDepth is more limited than with other datasets, and in particular with respect to IMC-PT. A possible justification is that the network is able to learn camera intrinsic priors of the MegaDepth images, which have less variability as outdoor scenes than IMC-PT. As discussed in the evaluation setup, all deep filters have been trained on SIFT, including FC-GNN, suggesting a less generalization to different image features with respect to handcrafted approaches based on geometric constraints like MOP.
Finally, unlike SIFT, directly applying NCC after the base pipeline without MOP provides some improvements, even if these are inferior to when NCC is preceded by MOP-base filtering. This confirming the original assumption that affine patch normalization is better than normalization by similarity for the refinement, but not as good as planar patch normalization.
@@ 4.2.4 SuperPoint+LightGlue
Table 3 reports the results for SuperPoint. Overall, the relative rank among the datasets and pipelines for SuperPoint is comparable with Key.Net. This is not accidental since both Key.Net and SuperPoint are based on corner-like keypoints, so the same considerations and observations done for Key.Net hold for SuperPoint. Moreover, in terms of absolute score among all the evaluated sparse or semi-dense pipelines, SuperPoint with MOP+MiHo+NCC achieves on MegaDepth the best AUC value, or a value almost equal to the best absolute AUC provided by LoFTR for ScanNet. The improvement in adding NCC in terms of recall and precision with respect to Key.Net is inferior but still relevant, implying a more accurate corner relocalization of the base SuperPoint.
Note that for the planar case, MOP+MiHo+ NCC is slightly inferior with respect to the base pipeline, achieving the top score, and with respect to MOP+MiHo, for which the score is almost the same of the base pipeline. It is reasonable considering that the base SuperPoint training procedure involves the homographic adaptation which is actually achieved by performing image matching on planar scenes, so that the base pipeline training overfits on this specific task.
MOP-based filters without NCC are better in fundamental matrix estimation when no MAGSAC is employed, while the corresponding MOP-based filters with NCC work better when MAGSAC follows. Figure 5(a) shows the distribution of the epipolar error of the matches as the next step of the pipeline is executed. NCC succeeds in refining most of the matches within a mid-range epipolar error, which for the specific configuration is about within 1 px and 5 px, as highlighted in the corresponding hollow in Fig 5(b). Nevertheless, NCC may fail in some cases as indicated by the small increase of the right tail of the plot, leading to higher errors for these few matches. Without MAGSAC the fundamental matrix estimation is not robust and hence a few strong outliers with high errors may affect the whole estimation. By adding MAGSAC these outliers are sweep out and the increased precision of the correctly refined matches improves the final result.
Finally, notice that results with Keypt2Subpx, which only refines and does not filter matches, are also included for MegaDepth and ScanNet with SuperPoint. While Keypt2Subpx is able to improve the results of the base pipeline, its performances are inferior to those of MOP+MiHo+NCC and FC-GNN. This highlights that also RANSAC-based global filters like MAGSAC can benefits from a local filtering stage.
@@ 4.2.5 LoFTR
Table 4 reports the results for LoFTR. Unlike other architectures relying on SuperGlue, MAGSAC greatly improves the LoFTR base pipeline. FC-GNN and MOP+MiHo+NCC provide the best results with the exception of ScanNet. Moreover, NCC on this latter dataset degrades the accuracy noticeably. These results are reasonable considering that LoFTR is a semi-dense architecture specifically designed for ScanNet scenes, and there are many keypoints corresponding to flat regions not feasible for NCC template matching.
@@
Excluding ScanNet, for the other datasets NCC greatly improves the AUCs, achieving the top scores, which are also the best as absolute values with respect to the sparse pipelines. This is highlighted by the high precision and the recall values exceeding those of the base LoFTR when using MOP-based filtering with NCC and without MAGSAC. As for previous base pipelines, FC-GNN is better than MOP-based filters with NCC in the case of pose estimation from the fundamental matrix.
4.2.6 Other sparse pipelines
Table 5 reports the results for ALIKED, while those for DISK and DeDoDe v2 can be found in Appendix E.3 since similar considerations hold. For ALIKED, the base pipeline is the best in almost all cases or very close to the best one, since by design ALIKED is trained to select only accurate matches and discard others. There are no match to be filtered and refined so that in the optimal case a filter or a refining module should not degrade the original score. Similar results on a different setups are reported [keypt2subpx]. In this sense, only MOP-based filters, AdaLAM and FC-GNN are robust and stable. Notice that FC-GNN only surpasses the base ALIKED on MegaDepth where it was trained on, and NCC can slight degrade MOP-based filters since the ALIKED matches are already sub-pixel accurate.
In the planar dataset, not involved in the ALIKED training, only MiHo-based filtering is better than the base pipeline while most of the other filters are slightly inferior to the base pipeline. The original ALIKED provides overall only robust and accurate base input matches differently from SuperPoint, that leads to better absolute AUC scores considering the base pipelines only. As a downside, this strict accuracy of the input limits the number of less probable correct matches which could be adjusted by a robust filtering approach, so that Superpoint matches filtered by MOP-based methods or FC-GNN provide overall remarkably better results.
Analogous observations hold for DISK, with the additional note that on IMC-PT the DISK base pipeline is much better than the corresponding base pipeline of SuperPoint, yet with FC-GNN or MOP-based filtering SuperPoint is able to almost fill the gap with DISK.
For DeDoDe v2 the results are quite unstable. Overall, this pipeline can only slight benefits by filtering and refinements, yet only using FC-GNN or MOP-based filters, but anyways the original or processed matches provide inferior results than those of SuperPoint.
4.2.7 Recent dense pipelines
Further analyses were done using the recent dense pipelines RoMa and MASt3R. While these pipelines provide SOTA accuracy with respect to the previous generation of the sparse approaches, these are currently less scalable and not suitable for general and common workflow tasks. The evaluations have been restricted to MegaDepth, ScanNet and the planar scenes, including MOP-based filtering, FC-GNN and Progressive-X. Detailed tables are reported in Appendix E.4.
For RoMa, filtering does not alter the accuracy, while NCC refinement strongly decreases the performances, likewise ALIKED. RoMa accuracy is generally higher than that of sparse pipelines and the RoMa dense approach provides many accurate matches. Note that MOP-based filters in most cases remove less matches than others in non-planar scenes. As for ALIKED, the accuracy of matches is very high, but since RoMa is dense there would be many keypoints on flat areas. Both these factors negatively contribute to NCC suitability, leading to significatively degraded performances.
MASt3R performances are better than RoMa on ScanNet and close to those of SuperPoint with filtering for MegaDepth. In any case, MOP-based filtering without NCC does not offer relevant benefits on these datasets. For the planar dataset, MASt3R performs very well with MOP-based filtering only with NCC, with an accuracy level comparable with that of RoMa and SuperPoint. Nevertheless, on non-planar scenes NCC strongly reduces the accuracy while the precision, and especially the recall, are oddly boosted. As presented in the next section through visual inspection, the accuracy of MASt3R matches is generally low due to its regular grid sampling strategy, but these inaccurate matches are consistent with the scene epipolar geometry. The MASt3R approach, inherited from DUSt3R, obtains the 3D structure of the scene for each image, aligns these 3D structures and then computes the outputs matches, which will be consistent with the 3D alignment and so will be the derived epipolar geometry. Please refer to the next visual analysis.